Hello,
I have a question I need help answering. Your time is appreciated.
How many liters of pure baking soda must be added to 200 liters of 30% baking soda solution to get 60% baking soda solution?
Thank you

Answer by mananth(16946)

------ percent ---------------- quantity
Baking Soda 30 ----------------200 L
Baking Soda 100 ----------------x L
Total 60 ----------------200+x L

30*200+100*x=60(200+x)

6000+100x=12000+60x

100 x - 60 x = 12000 - 6000

40 x = 6000

/ 40

x = 150 L 100 % Baking Soda